How do you solve for x
log 2x + log(x-5)=2?

Answer by scott8148(6628)

2x(x - 5) = 100

2x^2 - 10x - 100 = 0

x^2 - 5x - 50 = 0

(x + 5)(x - 10) = 0


remember, logarithms are not defined for values ≤0